OpenStudy (anonymous):

is F a continuous function of r? F(r) = gmr/R^3 if ror equal to R

OpenStudy (nikvist):

\[F(R-)=mg\frac{R}{R^3}=mg\frac{1}{R^2}\] \[F(R+)=mg\frac{1}{R^2}\quad\Rightarrow\quad F(R-)=F(R+)=F(R)\] F(r) is continuous